Project

General

Profile

NDOS Commissioning Meeting 2012/03/05

Attendance: Jon, Jan, Susan, Leon, Minerba, Martin, Denis, Peter, Mat,

Agenda:

  1. Beam and Detector Update
  2. DataCheck Report
  3. Installation Plans for the week
  4. DAQ Update and Plans for the week
  5. DCS Update and Plans for the week
  6. AOB

Beam and Detector Report

  • Peter: NuMI trigger timing showed a large just (~us) on Friday. The problem went away Sat. morning, but it looks like it came back today. Andrew is investigating. Peter thinks this looks like a difference between T-clock vs. B-synch; perhaps we are automatically switching over? Should know more by tomorrow.
  • Denis: we have not been producing MB spill files for the past month or so because MB has been upgrading their machines. He thinks that work should be complete, but needs to check.
  • Peter: we think the memory leak that was causing crashes in the event display has been fixed but new version of ART is required. Hopefully Andrew will be able to get to this later today. For now, turning off color shading will avoid the crash.

DataCheck Report (Jan):

  • Focused on consistent run period since 10/29/2011, where detector configuration has been quite stable: number of NuMI candidates per POT still shows a very large spread compared to the error bars. Jan required all subruns to have > 4000 active channels and to have lasted more than 10 minutes. See candppot.pdf and candppot_4200.pdf

Installation plans for this week

  • Leon: no plans for this week. Next week, we will try to install ~30 APDs on the detector. Jon, Mat and Leon will work on an installation scheme.

DAQ Update and Plans for next week

  • Peter: would like to try out multi-casting at some point. Kurt needs to write some configuration files, and some other minor updates will be necessary, so probably later this week.

DCS Update and Plans for next week

  • No report

AOB

  • None